LASER EXCITATION SPECTROSCOPY OF ^{58}NiH IN A MAGNETIC FIELD

\underline{CYRIL RICHARD}, HEATHER HARKER, PATRICK CROZET and AMANDA J. ROSS, LASIM, Université Lyon 1 & CNRS, 43 Bd du 11 novembre 1918, F-69622 Villeurbanne, France; DENNIS TOKARYK, Department of Physics and Center for Laser, Atomic, and Molecular Sciences, University of New Brunswick, Fredericton, Canada E3B 5A3.

Recent laboratory measurements of isotopologue ^{58}NiH by laser excitation around 17000−18000 cm^{−1} in a magnetic field have allowed us to study several electronic systems of this molecule. Zeeman patterns were analysed using literature values for ground state Landé factors^{a}. Effective electronic Landé factors geff have been determined. They show strong variation with parity in the rotational levels in some Ω′=3/2 states, giving evidence for extensive mixing between excited electronic states.

Effective electronic Landé factors geff

E(Ω = 3/2), v = 1 I(Ω = 3/2), v = 0
J  e  f  e  f
1.5 1.058  1.142  1.640  1.638
2.5 0.992  1.365  1.937  1.895
3.5 0.844  1.737  2.371  2.294
4.5 0.587  2.238  2.870  2.734
5.5 0.415  2.933  3.552  3.328
6.5 0.233  3.671  4.326  4.110

aMcCarthy et al. JCP \underline{107} (1997) 4179